OpenStudy (anonymous):

75.39 + b + 41.1=-28.946

OpenStudy (anonymous):

A/ 87.544 B/ 5.344 C/ -145.436 D/ -63.236

OpenStudy (anonymous):

Taking everything to the right hand side, b = -28.946 - 75.39 -41.1 = -(75.39+41.1+28.946) = -145.436 (C)

HanAkoSolo (jamierox4ev3r):

are you trying to solve for b? in this equation?

OpenStudy (anonymous):

That you so much, i really appreciate your help :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

No problem @AdrianGolfWang

OpenStudy (anonymous):

They solved it for me already Jamie but thanks anyway :)

HanAkoSolo (jamierox4ev3r):

al right :P as long as you understand it, that's good with me